OpenStudy (anonymous):

Establish the identity (sec x+tan x)(sec x-tanx)=1

OpenStudy (raden):

first, we can use the difference squares formula : (a+b)(a-b) = a^2 - b^2

OpenStudy (raden):

so, (sec x+tan x)(sec x-tanx) = sec^2 (x) - tan^2 (x), right ?

OpenStudy (raden):

then use the identity : sec^2 (x) = tan^2 (x) + 1

OpenStudy (raden):

therefore, (sec x+tan x)(sec x-tanx) = sec^2 (x) - tan^2 (x) = tan^2 (x) + 1 - tan^2 (x) = 1

OpenStudy (anonymous):

thank you =)

OpenStudy (raden):

you're welcome :)